OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 1, -8, 64, -512, …

OpenStudy (anonymous):

Again, the same forumla I said last time... http://en.wikipedia.org/wiki/Geometric_sequence 64/-8 = -8

OpenStudy (anonymous):

these are the answers that are shown???? an=-1(8)n − 1 an=8(-1)n − 1 an=1(-8)n − 1 an=-8(1)n − 1

OpenStudy (anonymous):

ah, sorry, I've just solved enough geomentric sequences that were asking for the common ratio that I didn't read carefully enough. I'm guessing that the n-1 on each of those is a subscript? If that's true then you should start at 1 because that's where your sequence starts. Then comes your common ratio (-8), and then you have the n-1, which is ALWAYS a way to show how many terms you wanted to solve for (in this case there are 4 terms, but the ... indicates that it goes on infinitely). \[1(-8)_{n-1}\]

OpenStudy (anonymous):

don't forget to medal